OpenStudy (anonymous):

The table below shows the velocity y, in miles per minute, of a toy car at different times x, in minutes: Time (x) (minutes) | 10 | 20 | 30| 40 ------------------------------------ Velocity (y) (miles) | 0.2| 0.8 | 0.1| 0.4 Part A: What is the most likely value of the correlation coefficient of the data in the table? Based on the correlation coefficient, describe the relationship between time and velocity of the toy car. [Choose the value of correlation coefficient from -1, 1, -0.04, 0.99]

OpenStudy (anonymous):

Part B: What is the value of the slope of the graph of velocity versus time between 10 minutes and 20 minutes, and what does the slope represent? Part C: Does the data in the table represent correlation or causation? Explain your answer.

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

I gtg to bed, @tester97 is beast ~

OpenStudy (the_fizicx99):

At Math*

OpenStudy (anonymous):

thanks anywat i know its late -.-

OpenStudy (anonymous):

anyway*

OpenStudy (the_fizicx99):

Yeah! This is the Part III for FLVS alg I xD I remember it, I haven't seen this in a while

OpenStudy (anonymous):

lol yesssss this is the last question and im DONE!

OpenStudy (tester97):

actually no that was all wrong

OpenStudy (tester97):

hold on

OpenStudy (anonymous):

ok

OpenStudy (tester97):

This all I got maybe @dan815 can help B. (y2 - y1)/(x2 - x1) = (0.8 - 0.2)/(20 - 10) = 0.6/10 = 0.06 the slope indicates that the velocity is increasing at a relatively slow rate from 10 mins to 20 mins. C. Zero correlation

OpenStudy (anonymous):

i appreciate all your help though!

OpenStudy (tester97):

Welcome :)

OpenStudy (dan815):

http://mathbits.com/MathBits/TISection/Statistics2/correlation.htm correlation coeff formula is there use it

OpenStudy (anonymous):

woahh that formula looks confusing..

OpenStudy (anonymous):

im getting no where trying to figure this out on my own

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ikram002p):

fill this table :-